# The quest for the GRAph Level autoEncoder (GRALE)

### v1

@@ A.3 Graph reconstruction metrics
To measure the error between a predicted graph $\hat{G}$ and a target graph $G$ we first consider the graph edit distance Gao et al., (2010); Sanfeliu and Fu, (1983), that is the number of modification (editions) that should be applied to get to $\hat{G}$ from $G$ (and vice-versa). The possible editions are node or edge addition, deletion, or modification, where node/edge modification stands for changing the label of a node/edge. In this paper, we set the cost of all editions to 1. It is well known that computing the edit distance is equivalent to solving a graph matching problem Aflalo et al., (2015). For instance, in the case of two graphs of the same size $G=(F,C)$ and $\hat{G}=(\hat{F},\hat{C})$, the graph edit distance is written as
$\texttt{Edit}(G,\hat{G})=\min_{P\in\sigma_{N}}\mathcal{L}_{\text{ALIGN}}(G,P[\hat{G}]),$ (12)
@@ B.1 Training Complexity
For the reconstruction accuracy, we observe overfitting for all datasets smaller than one million molecules. More interestingly, we observe that the performance on the downstream task is also highly dependent on the size of the pretraining dataset, as the performances keep improving past one million samples. Overall, it appears that this version of GRALE is able to leverage a very large molecular dataset. To apply GRALE to a different field, where pretraining datasets are smaller, it might require considering a different set of parameters of perhaps a more frugal architecture. For instance, the encoder and decoder could be replaced by the more lightweight baselines proposed in E.
@@ B.3 Reconstruction failure case
Appendix C GRALE architecture details
C.1 Featurizer
@@ C.2 Encoder
where $F^{1}\in\mathbb{R}^{n\times d_{F}}$ (resp. $C^{1}\in\mathbb{R}^{n\times n\times d_{c}}$) are initialized by applying a node-wise (resp. edge-wise) linear layer on $F$ (resp. $C$). The Evoformer Encoder layer used in GRALE is represented in Figure 13. Compared to the original implementation, we make two notable changes that make this version much more lightweight. First, we replace the MLP of the Feed Forward Block (FFB) with a simple linear layer followed by an activation function. Then, of the 4 modules dedicated to the update of $C$, we keep only one that we call Triangular Self Attention Block (tSAB) to highlight the symmetry with a Transformer Encoder. The definitions for all these blocks are provided in Appendix D.
@@
$X_{i}=\texttt{Linear}(F^{L}_{i})\quad\text{if}\quad i<n,\quad X_{i}=u\quad\text{otherwise.}$. (20)
@@
C.3 Decoder
The decoder takes as input the graph embedding $Z$ and should output both the reconstruction $\hat{G}=(\hat{h},\hat{F},\hat{C})$ and the node embeddings $\hat{X}$ that are required to match input and output graphs. To this end, the proposed architecture mimics that of a Transformer Encoder-Decoder except that we replace the usual Transformer Decoder with a novel Evoformer Decoder. More formally, the latent representation $Z=Z^{0}$ is first updated by a Transformer Encoder with $L$ layers
@@ C.3 Decoder
Figure 16: Similarly to a Transformer Decoder, the proposed Evoformer Decoder layer augments the Evoformer Encoder with 2 Cross-Attention Blocks (CAB) so that the output can be conditioned on some source $Z$. All the inner blocks are defined in D.
@@ C.5 Loss
Appendix D Definitions of attention based models inner blocks
@@
D.1 Node level blocks
@@
D.2 Graph level blocks
@@
Appendix E Ablation studies details
@@
Finally, we replace our loss by $\mathcal{L}_{\texttt{PIGVAE+}}(G,\hat{G},\hat{T})+\lambda\Omega_{\texttt{PIGVAE+}}(\hat{T})$ and we report the results for $\lambda=10$ (after a basic grid-search $\lambda\in\{0.1,1,10\}$).
@@
As detailed in C, the proposed featurizer $\phi$ augments the graph representation with high-order properties such as the shortest path matrix. We check the importance of this preprocessing step by removing it entirely. More precisely, we change the equation (16) that defines the node features into
@@
Appendix F Proofs of the theoretical results
F.1 Loss properties
@@
F.2 Positioning with respect to PIGVAE
In the following, we assume that the ground losses $\ell_{F}$ and $\ell_{C}$ are Bregman divergences as defined above. $G=(F,C)$ and $\hat{G}=(\hat{F},\hat{C})$ are graphs of size $N$ and we omit the padding vectors, enabling fair comparison with PIGVAE. In this context, the proposed loss is rewritten as
@@ F.2 Positioning with respect to PIGVAE
$\mathcal{L}_{\text{PIGVAE}}(G,\hat{G},\hat{T})=\mathcal{L}_{\text{ALIGN}}(G,\hat{T}[\hat{G}]),$ (57)

### v3

@@ A.3 Graph reconstruction metrics
[…]
@@ B.1 Training Complexity
For the reconstruction accuracy, we observe overfitting for all datasets smaller than one million molecules. More interestingly, we observe that the performance on the downstream task is also highly dependent on the size of the pretraining dataset, as the performances keep improving past one million samples. Overall, it appears that this version of GRALE is able to leverage a very large molecular dataset. Note that we also explore how a more frugal version of GRALE can be trained on smaller datasets in B.4.
@@ B.3 Reconstruction failure case
B.4 Training GRALE on small scale datasets
This paper is mostly devoted to the scenario where a large pretraining dataset is available, hence the focus on large, data-hungry, attention-based architectures. Yet, it is important to point out that GRALE can be adapted to smaller datasets by considering a more frugal architecture.
To illustate this we consider the BN dataset, a relatively small dataset of 200,000 8-node Bayesian networks Zhang et al., 2019a . Each network has a BIC score (on the Asia dataset) used as a regression target. To adapt GRALE to this setting, we replace the encoder and decoder by the more lightweight baselines proposed in F. We pre-trained GRALE on BN in an unsupervised fashion and used the learned graph-level embeddings as input to an MLP for the downstream regression task. We used the same hyperparameters as those reported for COLORING in table 9. We observe that GRALE outperforms all baselines by a significant margin, including AutoEncoders specifically designed for this dataset.
Table 11: Downsteam performances on the BN dataset. The reported baselines are taken from Zhang et al., 2019a .
Model RMSE $\downarrow$ Pearson’s $r$ ($\uparrow$)
GraphRNN $0.77$ $0.64$
GCN $0.55$ $0.83$
DeepGMG $0.78$ $0.62$
S-VAE $0.36$ $0.93$
D-VAE $\underline{0.30}$ $\underline{0.95}$
GRALE $\mathbf{0.18}$ $\mathbf{0.98}$
Appendix C Memory and Computational Complexity of GRALE
The theoretical complexity of GRALE with respect to the number of nodes $N$ is $\mathcal{O}(N^{3})$ in both time and memory. This stems directly from GRALE’s design: the graph attention layers (evoformer encoder and decoder), the Hungarian matching used in the matcher, and the graph reconstruction loss are all cubic. In this paper, this complexity motivated us to restrict our experiments to graphs of size $N\leq 32$, although this is not a strict limit.
We now report the practical cost (in time and memory) of training GRALE on larger graphs. Specifically, we measure the cost of running a single gradient step with batch size 1 on an RTX 3500 Ada (12GB VRAM), with all model hyperparameters set to the values reported in Table 9 for PUBCHEM32. The results are presented in Table 12. For example, training GRALE on graphs of size $N=128$ requires 3GB of GPU memory per graph in the batch. This suggests that 8 A40 GPUs (48GB each) could support training with a batch size of 128 and $N=128$, which is a reasonable configuration.
Table 12: GRALE computational cost per gradient step (batch size 1) on RTX 3500 Ada (12GB VRAM).
NN 8 16 32 64 128 256
Step Time (s) 0.01 0.01 0.02 0.02 0.03 0.07
GPU Memory (GB) 0.2 0.3 0.4 0.8 3.0 10.7
Appendix D GRALE architecture details
D.1 Featurizer
@@ D.2 Encoder
where $F^{1}\in\mathbb{R}^{n\times d_{F}}$ (resp. $C^{1}\in\mathbb{R}^{n\times n\times d_{c}}$) are initialized by applying a node-wise (resp. edge-wise) linear layer on $F$ (resp. $C$). The Evoformer Encoder layer used in GRALE is represented in Figure 12. Compared to the original implementation, we make two notable changes that make this version much more lightweight. First, we replace the MLP of the Feed Forward Block (FFB) with a simple linear layer followed by an activation function. Then, of the 4 modules dedicated to the update of $C$, we keep only one that we call Triangular Self Attention Block (tSAB) to highlight the symmetry with a Transformer Encoder. The definitions for all these blocks are provided in Appendix E.
@@
$X_{i}=\texttt{Linear}(F^{L}_{i})\quad\text{if}\quad i<n,\quad X_{i}=u\quad\text{otherwise.}$ (20)
@@
D.3 Decoder
[…]
@@ D.3 Decoder
Figure 15: Similarly to a Transformer Decoder, the proposed Evoformer Decoder layer augments the Evoformer Encoder with 2 Cross-Attention Blocks (CAB) so that the output can be conditioned on some source $Z$. All the inner blocks are defined in E.
@@ D.5 Loss
Appendix E Definitions of attention based models inner blocks
@@
E.1 Node level blocks
@@
E.2 Graph level blocks
@@
Appendix F Ablation studies details
@@
[…]
@@
As detailed in D, the proposed featurizer $\phi$ augments the graph representation with high-order properties such as the shortest path matrix. We check the importance of this preprocessing step by removing it entirely. More precisely, we change the equation (16) that defines the node features into
@@
Appendix G Proofs of the theoretical results
G.1 Loss properties
@@
G.2 Positioning with respect to PIGVAE
[…]
@@ G.2 Positioning with respect to PIGVAE
[…]